In a calorimeter, 50.0mL of AgNO3 and 50.0mL of HCl are mixed to yield 2.15g AgCl (molar mass 143.3g/mol) in the following reaction: Ag+(aq) + Cl-(aq) → AgCl(s)
The two solutions were originally at 23.0oC and the final temperature is 25oC. Calculate the heat of the reaction in kJ/mol of AgCl formed. Assume the final solution mass is 100.0g and the specific heat is that of water
1
Expert's answer
2020-05-12T11:34:49-0400

"\Delta"H = Q/"\nu" = c * m(H2O) * "\Delta"T / (m / M) = c * m(H2O) * M * "\Delta"T / m =

= 4.2 J / (g * oC) * 100 g * (25 - 23) oC * 143.3 g/mol / 2.15g = 55987 J/mol = 55.987 kJ/mol (approximately 56kJ/mol)

Answer: the heat of the reaction is 56kJ/mol